The base of a triangular sail is 15m shorter than its height.The area is 92m^2. Find the base and height of the sail.
Answer by sudhanshu_kmr(1152)


let height = x m ,base = x -15
1/2 * x * (x-15) = 92
=> x(x-15) = 184
=> x^2 - 15x -184 =0
=> x^2 -23x + 8x -184 =0
=> (x-23)(x+8) = 0
so, x =23
height = 23 , base = 8 m